# The Gradient of Mean Molecular Weight Across the Radius Valley

### v1

@@ 3.4 Population study of hybrid atmospheres
In Figure 7, we show calculations of the mean molecular weight across radius and orbital period. For what we term “reduced conditions”, $\log{f_{\rm O_{2}}}$ and $\log{f_{\rm S_{2}}}$ are uniformly sampled between $-11$ and $-1$ (cgs units; corresponding to $10^{-17}$ to $10^{-7}$ bar). When they are sampled uniformly between $-1$ to $2$ (cgs units; corresponding to $10^{-7}$ to $10^{-4}$ bar), “oxidized conditions” prevail. These ranges of values are arbitrary and only serve to illustrate that the gradient of $\mu$ across radius and orbital period is robustly present.
@@
[…]
Figure 9 comes with two caveats. Firstly, about $9\%$ of the outgassing calculations failed to attain numerical convergence when C/H is 10$\gtrsim$ 10$\times$ 10 × solar, presumably because the initial guesses for the volume mixing ratios of molecules are inadequate (see Appendix C). This implies that some of the empty regions in the $\mu$ versus C/H plot may actually be populated if all of the numerical calculations had converged. Secondly, carbon-rich atmospheres may precipitate graphite (Moses et al., 2013), a process we do not include in our CHOSSi chemical system. Graphite formation would remove carbon from the gaseous phase and reduce the mean molecular weight of the atmosphere.
@@ 4.3 What about miscible cores?
[…]
@@ 4.4.1 Activities, fugacities and solubilities
Geochemical outgassing calculations of super Earths and sub-Neptunes are limited mostly by the scarcity or non-existence of thermodynamic quantities needed to quantify non-ideal-gas behavior (fugacities), non-ideal mixing of gaseous components (activities) and the tendency for certain gaseous species to dissolve in melt (solubilities). While fugacities are not necessary for super Earths with surface pressures below 1 kbar, they are needed for sub-Neptune surface pressures $\sim 10$ kbar.
@@ 4.4.4 Large-scale transport in sub-Neptunian atmospheres
KH developed the theoretical formalism, derived the equations, performed the numerical calculations, made the figures and led the writing of the manuscript. MT co-developed the theoretical formalism, checked the mathematical derivations of KH and formulated both the equations and the computer code for calculating fugacity coefficients. JEO provided guidance on radius valley constraints and unpublished calculations from Rogers & Owen (2021). KH thanks Fabrice Gaillard for clarifying the solubility law of molecular hydrogen published in Gaillard et al. (2022), Janine Birnbaum for pointing out the tension between Dalton’s law and Newton’s second law, Nicolas Sartor for suggesting references on solubility laws, and Steve Mojzsis for stimulating conversations. KH acknowledges partial financial support from the European Research Council (ERC) Geoastronomy Synergy Grant (grant number 101166936).
@@ Appendix A Gibbs free energies of formation
[…]
@@
1 $-2.81990035\times 10^{2}$ $8.63857116\times 10^{-2}$ $1.14004268\times 10^{-7}$ $-8.64090135\times 10^{-11}$ $4.07761984\times 10^{-15}$
2 $-2.41212701\times 10^{2}$ $4.26423144\times 10^{-2}$ $3.75831753\times 10^{-6}$ $-3.67706178\times 10^{-10}$ $1.29138659\times 10^{-14}$
3 $-8.02022252e\times 10^{2}$ $1.54466780\times 10^{-3}$ $-2.87989051\times 10^{-7}$ $1.29417601\times 10^{-10}$ $-5.75526461\times 10^{-15}$
4 $-3.60405567\times 10^{2}$ $7.06878078\times 10^{-2}$ $5.98577039\times 10^{-7}$ $-7.12203913\times 10^{-11}$ $3.09693938\times 10^{-15}$
5 $-1.56860063\times 10^{2}$ $3.78126944\times 10^{-3}$ $8.70325866\times 10^{-7}$ $-5.98505785\times 10^{-11}$ $1.58640487\times 10^{-15}$
6 $-1.04479777\times 10^{2}$ $2.10409072\times 10^{-1}$ $7.36113799\times 10^{-6}$ $-9.39597816\times 10^{-10}$ $3.63793942\times 10^{-14}$
@@ B.2.1 Empirical coefficients for EoS of H2O
$a_{1}$ $-0.88517$dagger
$a_{2}$ $4.53\times 10^{-3}$dagger
$a_{3}$ $-1.3183\times 10^{-5}$dagger
$a_{4}$ $-0.22291$dagger
$a_{5}$ $-3.8022\times 10^{-4}$dagger
$a_{6}$ $1.7791\times 10^{-7}$dagger
@@ B.2.2 Empirical coefficients for EoS of CO2
$a_{1}$ $-0.10891$dagger
$a_{2}$ $-3.4203\times 10^{-4}$dagger
@@
$c_{0}$ $5.40776\times 10^{-3}$dagger
$c_{1}$ $-1.59046\times 10^{-6}$dagger
@@
$d_{0}$ $-1.78198\times 10^{-1}$dagger
$d_{1}$ $2.45317\times 10^{-5}$dagger
@@ B.2.3 Empirical coefficients for EoS of H2, CO and CH4
$\begin{split}a&=a_{0}\frac{T_{c}^{5/2}}{P_{c}}+a_{1}\frac{T_{c}^{3/2}T}{P_{c}},\\ b&=b_{0}\frac{T_{c}}{P_{c}},\\ c&=c_{0}\frac{T_{c}}{P_{c}^{3/2}}+c_{1}\frac{T}{P_{c}^{3/2}},\\ d&=d_{0}\frac{T_{c}}{P_{c}^{2}}+d_{1}\frac{T}{P_{c}^{2}},\\ \end{split}$ (B16)
@@ Appendix C Approximate CHOSSi system (ideal gas, ideal mixing, ignore solubility laws)
[…]
@@
[…]

### v2

@@ 3.4 Population study of hybrid atmospheres
In the top row of Figure 7, we show calculations of the mean molecular weight across radius and orbital period. For what we term “reduced conditions”, $\log{f_{\rm O_{2}}}$ and $\log{f_{\rm S_{2}}}$ are uniformly sampled between 11-11 and 1-1 (cgs units; corresponding to $10^{-17}$ to $10^{-7}$ bar). When they are sampled uniformly between 1-1 to $2$ (cgs units; corresponding to $10^{-7}$ to $10^{-4}$ bar), “oxidized conditions” prevail. These ranges of values are arbitrary and only serve to illustrate that the gradient of $\mu$ across radius and orbital period is robustly present. We note that random sampling of parameters to mitigate incomplete theoretical understanding has been successfully applied to the modeling of debris disks (Heng & Tremaine, 2010) and polluted white dwarfs (Veras & Heng, 2020).
In the bottom row of Figure 7, we show the corresponding values of the sampled oxygen fugacities relative to the iron-wüstite buffer ($\log{f_{\rm O_{2}}}-\mbox{IW}$). To compute the temperature and pressure dependence of IW, we use the empirical relation of Hirschmann (2021), which is strictly valid only up to 3000 K. In other words, extrapolation beyond this temperature limit is sometimes employed. The bottom-left panel of Figure 7 shows a qualitative trend that matches Figure 2c of Cherubim et al. (2025), who model an oxidation gradient across the radius valley by considering atmospheric escape. Quantitatively, both our sampled $\log{f_{\rm O_{2}}}$ (cgs units) and $\log{f_{\rm O_{2}}}-\mbox{IW}$ values are broadly consistent with those obtained by Cherubim et al. (2025) in their Figure 2c. Further comparison is unwarranted as the models of Cherubim et al. (2025) make multiple assumptions about atmospheric escape and planet formation history, including assuming an initial atmospheric abundance of oxygen that is solar and a core mass fraction of 0.3.
@@
The computed mean molecular weights in Figures 7 and 8 are insensitive to the carbon enrichment of the hydrogen envelope (sometimes termed the “metallicity”), which is assumed to be solar ($\mbox{C/H}=2.5\times 10^{-4}$). To prove this insensitivity, we execute an additional set of calculations where we allow $\log{\mbox{C/H}}$ to be uniformly sampled such that C/H varies from $0.1\times$ to $300\times$ its solar value. Figure 9 shows that there is no correlation between the upper envelope of $\mu$-values and C/H. However, when C/H is greater than about $20\times$ the solar value, it imposes a minimum value of $\mu$—exactly as one would expect for a “metal-rich” atmosphere. Despite sampling such a broad range of C/H values, the correlation between $\mu$ and $f_{\rm O_{2}}$ remains qualitatively identical: the upper envelope of mean molecular weight values depends on the oxygen fugacity. In other words, neither the gradient of $\mu$ nor its strength is primarily driven by the elemental abundance of carbon of the atmosphere. However, the range of $\mu$ values may be extended by high values of C/H. This is clearly seen when comparing the bottom left panel of Figure 8 with the right panel of Figure 9: for the former, the threshold value of the oxygen fugacity, for producing $\mu\gtrsim 2$ is $f_{\rm O_{2}}\gtrsim 10^{-11}$ bar; for the latter, there is no such threshold over the entire range of $f_{\rm O_{2}}$ sampled.
Figure 9 comes with two caveats. Firstly, about $8\%$ of the outgassing calculations failed to attain numerical convergence when C/H is $\gtrsim 10\times$ solar, presumably because the initial guesses for the volume mixing ratios of molecules are inadequate (see Appendix C). This implies that some of the empty regions in the $\mu$ versus C/H plot may actually be populated if all of the numerical calculations had converged. Secondly, carbon-rich atmospheres may precipitate graphite (Moses et al., 2013), a process we do not include in our CHOSSi chemical system. Graphite formation would remove carbon from the gaseous phase and reduce the mean molecular weight of the atmosphere.
@@ 4.3 What about miscible cores?
Benneke et al. (2024) have previously suggested that the enhanced value of the mean molecular weight inferred in the sub-Neptune TOI-270d may be attributed to it having a “miscible core”. While matter at pressures $\sim\mbox{GPa}$ certainly exists in a supercritical state, the main effect driving the enhanced mean molecular weight is the geochemical outgassing of these molten cores, from melts with temperatures of about 2000-4000 K, into the primordial hydrogen envelope. This conclusion is robust as it only depends on thermodynamics and Gibbs free energies, and occurs over a broad range of oxygen fugacities of the core.
@@ 4.4.1 Activities, fugacities and solubilities
Geochemical outgassing calculations of super Earths and sub-Neptunes are limited mostly by the scarcity or non-existence of thermodynamic quantities needed to quantify non-ideal-gas behavior (fugacities), non-ideal mixing of gaseous components (activities) and the tendency for certain gaseous species to dissolve in melt (solubilities). While fugacities are not necessary for super Earths with surface pressures below $\sim 1$ kbar, they are needed for sub-Neptune surface pressures $\sim 10$ kbar.
@@ 4.4.4 Large-scale transport in sub-Neptunian atmospheres
KH developed the theoretical formalism, derived the equations, performed the numerical calculations, made the figures and led the writing of the manuscript. MT co-developed the theoretical formalism, checked the mathematical derivations of KH and formulated both the equations and the computer code for calculating fugacity coefficients. JEO provided guidance on radius valley constraints and unpublished calculations from Rogers & Owen (2021). KH thanks Fabrice Gaillard for clarifying the solubility law of molecular hydrogen published in Gaillard et al. (2022), Janine Birnbaum for pointing out the tension between Dalton’s law and Newton’s second law, Nicolas Sartor for suggesting references on solubility laws, Steve Mojzsis for stimulating conversations and an anonymous referee for stimulating comments that improved the quality of the manuscript. KH acknowledges partial financial support from the European Research Council (ERC) Geoastronomy Synergy Grant (grant number 101166936; PIs: Gaillard, Heng and Mojzsis).
@@ Appendix A Gibbs free energies of formation
Figure 11 shows $\Delta G_{j}$ for j=1j=166 for the 6 net chemical reactions stated in equations (5) and (7). For convenience, we fit these $\Delta G_{j}$ curves using Chebyshev polynomials (denoted by ${\cal T}_{k}$ for a Chebychev polynomial of the $k$-th order),
@@
1 $2.81990035\times 10^{2}$ $8.63857116\times 10^{-2}$ $1.14004268\times 10^{-7}$ $8.64090135\times 10^{-11}$ $4.07761984\times 10^{-15}$
2 $2.41212701\times 10^{2}$ $4.26423144\times 10^{-2}$ $3.75831753\times 10^{-6}$ $3.67706178\times 10^{-10}$ $1.29138659\times 10^{-14}$
3 $8.02022252e\times 10^{2}$ $1.54466780\times 10^{-3}$ $2.87989051\times 10^{-7}$ $1.29417601\times 10^{-10}$ $5.75526461\times 10^{-15}$
4 $3.60405567\times 10^{2}$ $7.06878078\times 10^{-2}$ $5.98577039\times 10^{-7}$ $7.12203913\times 10^{-11}$ $3.09693938\times 10^{-15}$
5 $1.56860063\times 10^{2}$ $3.78126944\times 10^{-3}$ $8.70325866\times 10^{-7}$ $5.98505785\times 10^{-11}$ $1.58640487\times 10^{-15}$
6 $1.04479777\times 10^{2}$ $2.10409072\times 10^{-1}$ $7.36113799\times 10^{-6}$ $9.39597816\times 10^{-10}$ $3.63793942\times 10^{-14}$
@@ B.2.1 Empirical coefficients for EoS of H2O
$a_{1}$ 0.88517-0.88517 $\dagger$
$a_{2}$ $4.53\times 10^{-3}$ $\dagger$
$a_{3}$ $1.3183\times 10^{-5} \dagger$
$a_{4}$ 0.22291-0.22291 $\dagger$
$a_{5}$ $3.8022\times 10^{-4} \dagger$
$a_{6}$ $1.7791\times 10^{-7}$ $\dagger$
@@ B.2.2 Empirical coefficients for EoS of CO2
$a_{1}$ 0.10891-0.10891 $\dagger$
$a_{2}$ $3.4203\times 10^{-4} \dagger$
@@
$c_{0}$ $5.40776\times 10^{-3}$ $\dagger$
$c_{1}$ $1.59046\times 10^{-6} \dagger$
@@
$d_{0}$ $1.78198\times 10^{-1} \dagger$
$d_{1}$ $2.45317\times 10^{-5}$ $\dagger$
@@ B.2.3 Empirical coefficients for EoS of H2, CO and CH4
[…]
@@ Appendix C Approximate CHOSSi system (ideal gas, ideal mixing, ignore solubility laws)
where we have written $x\equiv\mbox{C/H}$ for compactness of notation. To lowest order, the surface pressure is independent of the elemental abundance of carbon (C/H) if $x\ll 1$.
@@
    where we have written $y\equiv\mbox{Si/O}$ and $F_{9}\equiv F_{8}/F_{2}$ for compactness of notation, even though it was not formally part of the system of equations.
@@
Appendix D Iron-wüstite (IW) Buffer
Figure 12: The oxygen fugacity (in bar) of the IW buffer using the empirical formulae of Ballhaus et al. (1991) and Hirschmann (2021), which are shown as dotted and solid curves, respectively.
The empirical formula of Ballhaus et al. (1991) is based on the experimental data of O’Neill (1987), which is valid for $833\leq T\leq 1450$ K. Ballhaus et al. (1991) used these data to produce an empirical fitting formula for IW (see their Table 2), which is restated in equation (27) of Tian & Heng (2024). Ballhaus et al. (1991) applied theoretical pressure corrections to derive this formula and its pressure range of validity is unclear.
Hirschmann (2021) extended the IW empirical formula to $1000\leq T\leq 3000$ K and pressures of $100\mbox{ kPa}\leq P\leq 100\mbox{ GPa}$. However, we have already noted that this does not prevent the oxygen fugacity from exceeding the total pressure under certain conditions.
Figure 12 shows the oxygen fugacity in absolute units (bar) of the IW buffer. The pressure dependence of IW is negligible between 1 bar (Earth-like surface pressures) and 10 kbar (sub-Neptune-like surface pressures). At pressures relevant for sub-Neptunes ($\lesssim 100$ kbar), the formulae of Ballhaus et al. (1991) and Hirschmann (2021) agree well (Figure 12). Unsurprisingly, these formulae diverge at $100\mbox{ GPa}=1$ Mbar, but such a pressure is irrelevant for sub-Neptunes. At this pressure, the oxygen fugacity decreases with temperature according to the empirical formulae.